If Jeff is 2 years younger than Gerald.
If the sum of their ages is 64
work out Jeffs age

Answer by ikleyn(52781)

.

Let x be the Jeff's age;  then the Gerald age is (x+2) years.


From the condition,

    x + (x+2) = 64.


It gives

    2x = 64 - 2 = 62

     x          = 62/2 = 31.


ANSWER.  Jeff is 31 years old.   Gerald is 31+2 = 33 years old.


CHECK.  The total age is  31 + 33 = 64 years.    ! Correct !

Solved.